# Hello Again! LLM-powered Personalized Agent for Long-term Dialogue

### v1

@@ 3.2 Results of Multi-Session Dialogue
We adopt two multi-session dialogue dataset to quantitatively evaluate our method in long-term dialogue scenarios. The first session is used to initialize conversation and the subsequent four sessions are used to evaluate the performance of long-term dialogue. In these experiments, our framework is applied to both zero-shot models, including ChatGLM and ChatGPT, and to tuned models such as BlenderBot and ChatGLM. The results are shown in Table 1.
@@
On both datasets, all models employing LD-Agent consistently achieve significant improvements across all sessions and metrics, showcasing the powerful ability of LD-Agent on supporting long-term dialogue. Most notably, comparing with previous state-of-the-art model HAHT, BlenderBot employing LD-Agent, which has similar parameter scale to HAHT, outperforms it with a large performance gap of 3.39%, 3.72%, 3.41%, and 3.32% on BLEU-2 ranging from session 2 to 5. This further highlighting the effectiveness of LD-Agent on long-term dialogue tasks.
@@
3.3 Ablation Studies
The results clearly demonstrate that all modules positively influence long-term dialogue capabilities, with the event memory module contributing the most significant improvements. It is worth noting that although all modules experience a performance decline as the number of sessions increased, the addition of the event memory module results in more stable performance compared to the use of user or agent personas. This highlights the critical role of event memory in maintaining coherence across multiple sessions.
@@ 3.4 Persona Extraction Analysis
To explore the effect of different persona extractor, including zero-shot ChatGLM with Chain-of-Thought [32] and ChatGLM tuned on the persona extraction dataset collected from MSC training set, we carry out comparison experiments on two perspectives: Persona Extraction Accuracy and Impact to Response Generation. The results are shown in Table 3.
@@
We evaluate the extraction accuracy on the persona extraction dataset collected from MSC testing set, through BLEU-2/3, R-L, and ACC. ACC is to assess the classification accuracy of dividing utterance into “with personas” or “without personas”. The results of extraction in Table 3 show that the extractor after tuning performs better than CoT ChatGLM on all metrics. The higher BLEU and R-L indicates the tuned extractor performs better capability to extract personas from an utterance, while higher ACC indicates a stronger capability to distinguish whether personas are contained in an utterance.
@@ 3.5 Human Evaluation
To further explore the performance of LD-Agent in real-life conversation, we adopt human evaluation to separately evaluate the ability of memory recall and response generation with the results on Figure 3.
@@
Retrieval mechanism plays a crucial role for event memory accurately utilized in long-term dialogue. To evaluate the superiority of topic-based retrieval approach than direct semantic retrieval commonly used in previous methods, we conduct an event memory human evaluation. In the beginning, we initialize a conversation using first four sessions and store event memories for each session into long-term memory bank. In the last session, we let evaluators select relevant memories from long-term memory bank for each utterance as the ground truths. Consequently, we separately utilize direct semantic retrieval and topic-based retrieval to search relevant memories for each utterance, and calculate the accuracy and recall based on human annotations. The results are shown in Figure 3a. The topic-based retrieval outperforms direct semantic retrieval with significant gap on both ACC and Recall, proving that our retrieval method accurately retrieves relevant memories.
(a) Comparison of Different Retrieval Mechanism
(b) Comparison of Response Generation
@@
To further validate the superiority of LD-Agent in long-term open-domain dialogue tasks, we organize multiple multi-session human-bot conversations on ChatGLM with LD-Agent and w/o LD-Agent. We first initialize a predefined dialogue as the first session for all chatbots. Subsequently, we employ some human evaluators to chat with each chatbot with a time interval from first session. The interactions are evaluated on three aspects: coherence, fluency and engagingness. The results in Figure 3b demonstrate the advantages of LD-Agent in long-term real-life dialogue scenarios.
@@
Table 4: The results of cross-domain evaluation on MSC and CC. “Zero-shot” indicates the ChatGLM without tuning. “CC-tuning” indicates the ChatGLM tuned on CC dataset. “MSC-tuning” indicates the ChatGLM tuned on MSC dataset.

### v2

@@ 4.3 Results of Multi-Session Dialogue
We adopt two multi-session dialogue dataset to evaluate our method in long-term dialogue scenarios. The first session is used to initialize conversation and the subsequent four sessions are used to evaluate the performance of long-term dialogue. To ensure consistency with real-world scenarios, we simulate all previous dialogues from the start before calculating each utterance’s generation accuracy. In these experiments, LD-Agent is applied to both zero-shot models, including ChatGLM and ChatGPT, and to tuned models like BlenderBot and ChatGLM with the results reported in Table 1.
@@
On both datasets, all models using LD-Agent consistently achieve significant improvements across all sessions and metrics, showcasing the powerful ability of LD-Agent on supporting long-term dialogue. Most notably, compared to previous state-of-the-art model HAHT, BlenderBot employing LD-Agent with similar parameter scale to HAHT, outperforms it with a large performance gap on BLEU-2 ranging from session 2 to 5 on both datasets. This further highlights the effectiveness of LD-Agent.
@@
4.4 Ablation Studies
The results clearly prove that all modules positively influence long-term dialogue capabilities, with the event memory module contributing the most. It is worth noting that although all modules experience a performance decline as sessions increase, the addition of the event memory module results in more stable performance compared to the use of user or agent personas. This highlights the critical role of event memory in maintaining coherence across multiple sessions.
@@ 4.5 Persona Extraction Analysis
To explore the effect of different persona extractor, including zero-shot ChatGLM with Chain-of-Thought Wei et al. (2022b) and ChatGLM tuned on the persona extraction dataset collected from MSC training set, we carry out comparison experiments on two perspectives: Persona Extraction Accuracy and Impact to Response Generation. The results are shown in Table 3.
@@
We evaluate the extraction accuracy on the persona extraction dataset collected from MSC testing set, through BLEU-2/3, R-L, and ACC. ACC is to assess the classification accuracy of dividing utterance into “with personas” or “without personas”. The results of extraction in Table 3 show that the extractor after tuning performs better than CoT on all metrics. The higher BLEU and R-L indicates the tuned extractor performs better capability to extract personas, while higher ACC indicates a stronger capability to distinguish whether personas are contained in an utterance.
@@ 4.6 Human Evaluation
To further explore the performance of LD-Agent in real-life conversation, we employ 8 students to assist with the assessment on memory recall and response generation according to the guidelines in Figure 4. More human evaluation details can be found in Appendix. 4.2.
@@
Retrieval mechanism plays a crucial role for event memory accurately utilized in long-term dialogue. To evaluate the superiority of topic-based retrieval than direct semantic retrieval commonly used in previous methods, we conduct an event memory human evaluation. We first initialize a conversation using first four sessions and store event memories for each session into long-term memory bank. In the last session, we let evaluators select relevant memories from long-term memory bank for each utterance as the ground truths. Consequently, we separately utilize direct semantic retrieval and topic-based retrieval to search relevant memories for each utterance, and calculate the accuracy and recall based on human annotations. The results are shown in Figure 3(a). The topic-based retrieval outperforms direct semantic retrieval with significant gap on both ACC and Recall, proving that our retrieval method accurately retrieves relevant memories.
@@
To further validate the superiority of LD-Agent in long-term open-domain dialogue tasks, we organize multiple multi-session human-bot conversations on ChatGLM with LD-Agent and w/o LD-Agent. We first initialize a predefined dialogue as the first session for all chatbots. Subsequently, we employ some human evaluators to chat with each chatbot with a time interval from first session. We ask human evaluators engage in 2-3 chat sessions with both the original ChatGLM and LD-Agent according to their own thoughts. Afterward, the interactions will be evaluated on three aspects: coherence, fluency and engagingness. The results in Figure 3(b) demonstrate the advantages of LD-Agent in long-term real-life dialogue scenarios.
Session 2 Session 3
Model BL-2 BL-3 R-L BL-2 BL-3 R-L
Zero-shot 9.53 4.82 22.76 9.22 4.43 22.18
Zero-shotLDA 8.94 4.44 21.54 8.34 4.03 21.00
MSC-tuning 8.37 3.88 22.93 8.49 3.99 22.96
MSC-tuningLDA 21.71 15.42 34.97 20.87 14.74 34.01
CC-tuning 15.89 9.90 30.59 15.97 10.06 30.27
CC-tuningLDA 25.69 19.53 39.67 25.93 19.72 39.15
Session 4 Session 5
Zero-shot 9.15 4.48 22.18 8.99 4.43 22.10
Zero-shotLDA 8.28 3.82 20.67 8.12 3.81 20.54
MSC-tuning 7.97 3.75 22.15 7.60 3.70 21.87
MSC-tuningLDA 19.57 13.51 32.72 18.59 12.80 31.68
CC-tuning 16.10 10.31 30.54 15.10 9.34 29.43
CC-tuningLDA 25.82 19.40 39.05 24.26 18.16 37.61
@@
Table 4: The results of cross-domain evaluation on CC. “Zero-shot” indicates the ChatGLM without tuning. “CC-tuning” indicates the ChatGLM tuned on CC. “MSC-tuning” indicates the ChatGLM tuned on MSC.
4.7 Generality Analysis
We further explore its generality from two perspectives: cross-domain and cross-task capability.
Cross-domain Results.
The cross-domain capability is crucial for open-domain dialogue task. Poor cross-domain performance, common in models tuned with specific datasets, limits their real-world practicality. To assess our tuned model’s real-world potential, we conduct cross-evaluation on MSC and CC, two datasets with significant domain gaps due to different collection methods, including manual annotation and LLM generation. We first tune ChatGLM on MSC and test it on CC, then reverse the process. We show the results on CC in Table 4, and the full results on MSC and CC in Table 8 of Appendix. It shows that models tuned on one dataset still performs well on the other dataset, only with a slight performance decrease than the models tuned on the same dataset. Besides, cross-domain tuned models consistently outperform zero-shot models by a significant margin. These experiments highlight strong cross-domain and practical potential of LD-Agent.
Model BL-1 BL-2 BL-3 BL-4 MET R-L
GPT-2 10.37 3.60 1.66 0.93 4.01 9.53
GSN 10.23 3.57 1.70 0.97 4.10 9.91
HeterMPCBART 12.26 4.80 2.42 1.49 4.94 11.20
BART 11.25 4.02 1.78 0.95 4.46 9.90
BARTLDA 14.40 4.92 2.07 1.00 5.30 12.28
Table 5: Multi-party performance on the Ubuntu IRC.
Cross-task Results.
The other capability worth exploring is the transferability of LD-Agent to different dialogue tasks. We explore the effectiveness of our method on multiparty dialogue, a task requires playing multiple roles simultaneously. We conduct experiments on Ubuntu IRC dataset Hu et al. (2019), a commonly used multiparty dialogue dataset. where our backbone adopts BART Lewis et al. (2020). We compare our method with some previous multiparty dialogue methods, including GPT-2 Radford et al. , GSN Hu et al. (2019), HeterMPCBART Gu et al. (2022), and BART tuned without prompt. The results are reported at Table 5. It can be seen that BART tuned with LD-Agent obtained the state-of-the-art performance in most metrics, outperforming previous multiparty dialogue approach HeterMPCBART, which also employs BART as backbone. This well proves the powerful task transferability of LD-Agent.
5 Conclusion
In this work, we delved into the long-term open-domain dialogue agent to meet the real-life chatbot demands for long-term companionship and personalized interactions. Unlike most prior solutions, which primarily focus on brief conversations spanning 2-15 turns within a single session, long-term dialogue agents could support consistent interactions over extended periods, even with significant time gaps between sessions. These agents can also perceive and adapt to user personalities, enabling them to deliver more accurate and personalized services. To achieve this, we introduced a general long-term dialogue agent framework, LD-Agent, which comprehensively considers both historical events and user-agent personas to support coherent and consistent conversation. Its decomposition into three learnable modules significantly enhances its adaptability and transferability. Extensive experiments demonstrated LD-Agent’s strong capability in long-term dialogue tasks, showing its practicality across multiple benchmarks, models, and tasks.
Limitations
Though LD-Agent exhibits impressive effectiveness and generality on long-term dialogue, we believe that the research on long-term open-domain dialogue still has a long way to go. For instance, there are some remained limitations of this work from the following perspectives:
Lacking Real-World Datasets.
Current long dialogue datasets are typically synthetic, created manually Xu et al. (2022a) or generated by large language models Jang et al. (2023); Maharana et al. (2024), which introduces a gap from real-world data. Due to the challenges in gathering authentic long-term dialogue data, our work is currently confined to these synthetic datasets. We aim to validate our approach on real long-term dialogue data in the future.
Sophisticated Module Design.
In this paper, LD-Agent provides a general framework for long-term dialogue that allows for modular optimization. However, the module implementations only employ some basic methods without more sophisticated design, which can be further explored in the future. For the memory module, long-term memory summarization Wang et al. (2023c) and accurate memory retrieval Gao et al. (2023b) are two critical techniques worth further exploration. For the persona module, improving methods for personality extraction Wu et al. (2020b) and persona-based retrieval Gu et al. (2021); Oh et al. (2023); Kasahara et al. (2022) offers promising directions for future work.
Acknowledgments
This research was supported by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (No. MSS24C004).
Appendix
In this Appendix, we discuss the following topics: (1): We elaborate more detailed experimental settings in Appendix. A. (2): We conduct various qualitative analysis in Appendix. B. (3): More experimental results are shown in Appendix. C. (4): In the Appendix. D, the prompts utilized in LD-Agent is illustrated.
Appendix A Detailed Evaluation Settings
A.1 Datasets
Multi-session Dataset.
Our experiments are conducted on two illustrative multi-session datasets: MSC Xu et al. (2022a) and CC Jang et al. (2023). Both datasets feature 5 sessions, with approximately 50 conversational turns per sample. Both of MSC and CC have the time interval annotations across sessions, which are employed to make the time decay factor work. MSC extends the PersonaChat dataset Zhang et al. (2018), utilizing PersonaChat for the initial session and employing human-human crowd workers to simulate the dialogues in subsequent sessions. The time intervals between sessions can span several days, and the dataset includes records of the participants’ personas. We follow the split of Zhang et al. (2022) with 4,000 conversations for training, 500 conversations for validation, and 501 conversations for testing. CC is complied by ChatGPT, which guides interactions according to a predefined event graph and participant relationships, with time intervals between sessions extending over several years. We employ the same data scale as MSC, with 4,000 conversations for training, 500 conversations for validation, and 501 conversations for testing.
Dialog Summary Dataset.
We utilize the DialogSum dataset Chen et al. (2021) for event summary. The dataset contains 13,460 dialogues with corresponding manually labeled summaries and topics. 12,460 dialogues used for training, 500 samples for validation, and 1,500 for test.
Persona Extraction Dataset.
We directly use the personas annotations from the MSC to construct the personas extraction dataset. The dataset is divided into train and valid sets, with the train set containing 26,605 samples and the valid set containing 17,660 samples.
Response Generation Dataset.
The response generation dataset is constructed based on the original conversations provided by MSC and CC, combined with the relevant memories and personas generated and extracted through our framework, and constructed using the prompt from Appendix. D.1. Among them, 34,907 samples are used for training, and 11,851 samples are used for validation.
Multi-party Dataset.
To explore the transferability of LD-Agent on other dialogue tasks. We apply our method to the Ubuntu IRC benchmark Hu et al. (2019), a dataset of multiparty tasks. We follow the split of previous works Hu et al. (2019); Gu et al. (2022) with 311,725 dialogues for training, 5,000 dialogues for validation, and 5,000 dialogues for testing.
A.2 Metrics
Automatic Evaluation Metrics.
BLEU-N Papineni et al. (2002) (BL-N) and ROUGE-L Lin (2004) (R-L) metrics are commonly used automatic evaluation metrics in dialogue generation tasks. BLEU-N measures N-gram overlaps between the generated text and the reference text, while ROUGE-L focuses on sequential coherence. We employ the METEOR (MET) Banerjee and Lavie (2005) metric in multi-party tasks as a complement to the BLEU metric, enhancing it with synonym calculation capabilities. In addition, accuracy (ACC) is calculated to measure the classification accuracy of different persona extractors. To further validate the diversity of the generated responses, we also employ the Distinct-1/2/3 (Dist-1/2/3) metrics for evaluation.
Human Evaluation Metrics.
The human evaluation consists of two parts: 1) retrieval mechanism evaluation, and 2) response generation evaluation. For the former, we use Accuracy (ACC) and Recall to measure the effectiveness of the memory retriever. Accuracy reflects the extent to which the retriever’s results align with those of a human retriever. Recall indicates the frequency with which the retriever correctly identifies the presence of relevant memories in the memory database (with an optimal Recall value of 100, as we evaluate the most recent session). For the response generation, we evaluate LD-Agent on three aspects: coherence, fluency, and engagingness. Coherence measures the chatbot’s capabilities to maintain the coherence of topic and logic across sessions. Fluency reflects the natural and fluent degree of interactions, making the interaction similar to human-human interactions. Engagingness measures a user’s interest in interacting with the target chatbot.
A.3 Baselines
  • HAHT Zhang et al. (2022): This is the state-of-the-art model crafted for multi-session, open-domain dialogue. It encodes all historical information and utilizes an attention mechanism to capture the relevant information to an ongoing conversation.
  • BlenderBot Roller et al. (2021): This is a commonly used large-scale open-domain dialogue model pre-trained on online social discussion data, who has the similar parameter scale with HAHT. We can achieve a fair comparison with HAHT by deploying LD-Agent on BlenderBot.
  • ChatGLM3 Zeng et al. (2023): This is an offline large language model 6B parameters. The model is pre-trained on 1T corpus, performing remarkable zero-shot reasoning capabilities.
  • BART Lewis et al. (2020): This is a denoising autoencoder with transformer architecture, trained to reconstruct original text from corrupting text.
A.4 Implementation Details
For the event summarizer, persona extractor, and response generator modules, we utilize the same base model for them and employ the LoRA mechanism across all configurations. All training and evaluation operated on a single NVIDIA A100 GPU. For the ChatGLM3-6B, it is optimized by an Adam Kingma and Ba (2015) optimizer with the learning rate of 5e-5. We configure this model with a batch size of 4 and train it over 3 epochs. For BlenderBot, the initial learning rate is set to 2e-5, with the batch size and the number of training epochs set at 4 and 5, respectively. Moreover, the interval time threshold $\beta$ is set to 600 seconds, while the temperature coefficient for the time decay coefficient $\tau$ is set to 1e+7, and the semantic cosine similarity threshold $\gamma$ is set to 0.5
(a) Retrieval Mechanism Guidelines
(b) Response Generation Guidelines
Figure 4: Human Evaluation Guidelines.
Appendix B Qualitative Analysis
We have conducted various quantitative experiments in our main paper. However, there is not a single "gold" reference answer in practical open-domain dialogue scenarios. To further verify the superiority of LD-Agent, we conduct some additional qualitative analysis in this section.
Figure 5: Illustration of summary module impact.
B.1 Persona Ablation
In Section 4.4, We conduct ablation studies to evaluate the role of the persona module. To explore its specific impact on dialogue process, we provide the same user query and observe the differences in LD-Agent’s responses with and without the initial personas. As shown in Figure 5, the response using the initial personas proactively inquires about the new company, aligning with the "enthusiastic" and "proactive" personas.
B.2 Memory Ablation
Moreover, we explore the specific impact of memory module on dialogue process. We first manually initialize a historical session and then compare the responses in a new session with and without the memory module. As shown in Figure 6, the Agent with the memory module effectively recalls the movies recommended in the historical session, effectively enhancing the continuity of the conversation.
Figure 6: Illustration of memory module impact.
B.3 Event Summarizer Analysis
In Section 3.2.1, we introduce how we train an event summarizer based on DialogSum to extract more concise event summaries. To validate the superiority of our trained summarizer than original ChatGLM, we explore it from three aspects: 1) in-domain (ID) evaluation; 2) out-of-distribution (OOD) evaluation; 3) module impact analysis.
We first evaluate the trained summarizer on two dialogue summarization test sets, DialogSum (Chen et al., 2021) and SAMSum (Gliwa et al., 2019). The former is consistent with the domain of the summarizer’s training set, while the latter is an out-of-domain dataset. The results in Table 6 show that the trained summarizer achieves consistent improvements over zero-shot ChatGLM across all metrics on both datasets. The improvements on the OOD dataset demonstrate the stronger generalization ability of our summarizer.
ID evaluation (DialogSum)
Model BL-1 BL-2 BL-3 BL-4 R-L Dist-1 Dist-2 Dist-3
ChatGLM 22.36 11.09 5.39 2.71 24.22 75.83 92.76 98.13
LD-Agent 40.71 24.31 15.55 9.40 40.31 87.53 98.74 99.67
OOD evaluation (SAMSum)
Model BL-1 BL-2 BL-3 BL-4 R-L Dist-1 Dist-2 Dist-3
ChatGLM 30.82 18.34 11.57 7.04 32.81 75.96 94.48 98.60
LD-Agent 33.98 19.73 12.38 7.42 37.28 93.05 99.40 99.79
Table 6: In-Domain and Out-of-Domain Evaluation
We then present in Figure 7 the summarization results of the same dialogue context generated by zero-shot ChatGLM and the LD-Agent summarizer. It can be seen that the summary generated by LD-Agent is more concise and effectively conveys the most important information in the conversation, making it more suitable for long-term dialogue given the growing demand for memory storage.
Figure 7: Illustration of event summary comparison.
Figure 8: Illustration of persona extraction comparison.
B.4 Persona Extractor Analysis
In Section 3.3, we introduce our trained dynamic personas extractor. To further explore its effectiveness, we conduct a comparison between our method and zero-shot ChatGLM for persona extraction. In Figure 8, we utilized an utterance from real-world scenarios as input and observe that the persona extracted by our tuned extractor is more concise and logical, making long-term dialogue processes more feasible.
B.5 Response Generation Analysis
To further analyze the generation ability of LD-Agent in long-term dialogue, we illustrate an example in Figure 10. It can be seen that the response generated by LD-Agent successfully captures the information about “General Nathan Bedford Forrest” they talked about in the history session, which performs better than original ChatGLM.
Appendix C Additional Experimental Results
In this section, we introduce some additional experimental results.
C.1 Part of Speech Importance Analysis
In Section 3.2.1, we implement a topic-based retrieval mechanism. Specifically, we compute the overlap of nouns to represent topic similarity. The rationale for using nouns to determine relevance lies in the fact that nouns typically convey more substantial information. We substantiate this by examining the concept of information entropy. Information entropy is often reflected through word frequency, with less frequent words generally carrying more information, thereby exhibiting higher information entropy. Figure 9 illustrates the average information entropy across various parts of speech, calculated from the DialogSum (Chen et al., 2021) dataset, showing that nouns possess the highest average information entropy.
Figure 9: Part of Speech Entropy Comparison.
C.2 Generation Diversity Analysis
To further validate the diversity of the generated responses by LD-Agent, we employ Dist-1/2/3 metrics to conduct evaluation on MSC. The results are shown in Table 7. We can observe that the responses generated by LD-Agent are consistently more diverse than generated by ChatGLM, indicating the powerful generation capability of our generator.
Session 2 Session 3
Model Dist-1 Dist-2 Dist-3 Dist-1 Dist-2 Dist-3
ChatGLM 81.71 92.51 95.54 79.32 90.79 94.18
LD-Agent 86.14 94.51 96.66 83.00 93.45 96.10
Session 4 Session 5
ChatGLM 78.41 90.13 93.68 78.07 89.88 93.46
LD-Agent 81.01 92.43 95.36 86.43 95.50 97.46
Table 7: The diversity of response generation on MSC.
Appendix D Prompt
D.1 Prompt of Event Summary
D.2 Prompt of Persona Extraction
D.3 Prompt of Response Generation
Prompt 4:  Agent Response Generation Prompt $\mathcal{SYS\ PROMPT}$: As a communication expert with outstanding communication habits, you embody the role of $\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\{agent\ name\}$ throughout the following dialogues. Here are some of your distinctive personal traits: $\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\{agent\ traits\}$.    $\mathcal{USER\ PROMPT}$: <CONTEXT> Drawing from your recent conversation with $\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\{user\ name\}$: $\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\{context\}$ <MEMORY> The memories linked to the ongoing conversation are: $\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\{memories\}$ <USER  TRAITS> During the conversation process between you and $\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\{user\ name\}$ in the past, you found that the $\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\{user\ name\}$ has the following characteristics: $\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\{user\ traits\}$ Now, please role-play as $\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\{agent\ name\}$ to continue the dialogue between $\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\{agent\ name\}$ and $\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\{user\ name\}$. $\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\{user\ name\}$ just said: $\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\{input\}$ Please respond to $\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\{user\ name\}$’s statement using the following format (maximum 30 words, must be in English): RESPONSE: